an open box is to be constructed from a square piece of sheet metal by removing a square of side 1 foot from each corner and turning up the edges. if the box is to hold 4 cubic feet, what should be the dimensions of the sheet metal?
some how I got that the dimension of the square should be 2ft by 2ft but that doesn't sound right at all.

Answer by Edwin McCravy(20055)

an open box is to be constructed from a square piece of sheet metal by removing a square of side 1 foot from each corner and turning up the edges. if the box is to hold 4 cubic feet, what should be the dimensions of the sheet metal?
We start with this square. Let the square have dimensions s ft × s ft



We partition it this way, the squares to cut out are 1 ft each, and
there are two from each side so the middle piece is s-2



Then we cut away the squares and have this:



When we fold up the flaps we will have a rectangular solid

whose length = L = s-2, whose width = W = s-2, and whose height = H = 1

V = LWH

and we are given that the volume of the box is 4 cubic feet, so

4 = (s-2)(s-2)(1)
4 = (s-2)(s-2)
4 = s²-2s-2s+4
4 = s²-4s+4
0 = s²-4s
0 = s(s-4)

s = 0  s-4 = 0
       s = 4

We ignore the zero solution.  So the sheet to cut 
should be 4 ft × 4 ft.

Edwin